OpenStudy (anonymous):

Use substitution to solve the system of equations. 2x+y=-7 -5x+2y=22 A. (–36, 65) B. (–4, 1) C. (8, –23) D. (36, –79)

OpenStudy (igreen):

Rearrange and substitute: \(2x + y = -7\) \(2x = -y - 7\) \(x = -1/2y - 7/2\) Substitute: \(-5x + 2y = 22\) \(-5(-1/2y - 7/2) + 2y = 22\) Distribute -5 into the parenthesis: \(2.5y + 17.5 + 2y = 22\) Subtract 17.5 from both sides: \(2.5y + 2y = 4.5\) Combine like terms: \(4.5y = 4.5\) Divide 4.5 to both sides: \(y = 1\)

OpenStudy (igreen):

Our \(y\) value is 1. And there's only one choice with the y-value of 1. So your answer is C.

OpenStudy (igreen):

So your answer is B*

OpenStudy (anonymous):

Thanks again. what about this one??

OpenStudy (anonymous):

Which of the following is not a solution to the system of linear equations below? x-2y=4 -2x+4y=-8 A. (2, –1) B. (4, 0) C. (0, 2) D. (6, 1)

OpenStudy (igreen):

C